OpenStudy (anonymous):

A payday loan company charges 4 percent interest for a two-week period. What would be the annual interest rate from that company?

jimthompson5910 (jim_thompson5910):

There are 52 weeks in a year, so 52/2 = 26 two-week periods in a year. You need to solve 4/1 = x/26 for x to get your answer

OpenStudy (anonymous):

4/26?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

cross multiply

OpenStudy (anonymous):

4*26?